# Learning to Collaborate: A Capability Vectors-based Architecture for Adaptive Human-AI Decision Making

### v1

@@ 4.2 Learning mechanism of capability vectors
To learn the capability vectors, we assign each expert an unique one-hot vector as the initial capability representation, and then a linear transformation is applied for convert the one-hot capability representation to capability vector. As the parameters of the linear transformation are learnable, then the obtained capability vectors can be considered as learnable ones (from scratch). We use Figure  2 to illustrate the workflow of this mechanism. We use a matrix comprised of the one-hot vectors of all or a subset of selected decision makers, and do matrix multiplication with a capability matrix of all decision makers to get the matrix of selected capability vectors. It is then concatenate with a separation embedding and a task embedding. The concatenated vector sequence is input to the weights generation model shown in Figure 1. The capability matrix can be updated with back-propagation through the gradients from the final decision errors. This mechanism is comparable with the learning of word/token embeddings for pretrained language models, while each decision maker can be considered as a token in the vocabulary. The difference is that the order of capability vectors does not matter for the default collaborative decision mechanism presented in Figure 1 (we believe that there will be alternative mechanisms where the order do matter, which can be considered in future works). Moreover, there is an extra task embedding that should be distinguished from the capability vector sequence. Usually, the size of capability embedding vectors $d_{c}$ should be much smaller than the vocabulary size $d_{v}$ for capturing the relationships between different decision makers (e.g. capability similarity or complementarity) and achieving better generalization. We provide two strategies: a. Use a transformer encoder weight generation model with a smaller hidden size, which may require an extra linear mapping for reducing the dimension of task embedding; b. Do a low-rank decomposition on the original capability matrix $\mathbf{C}\rightarrow\mathbf{U}\mathbf{V}^{T}$, where $\mathbf{C}\in\mathbb{R}^{d_{v}\times d_{e}},\mathbf{U}\in\mathbb{R}^{d_{v}\times d_{c}},\mathbf{V}\in\mathbb{R}^{d_{c}\times d_{e}}$ ($d_{c}<<d_{v}$) and consider the intermediate representation with lower dimension as the capability representation. We will compare the two strategies in the experiments.
@@ 4.3 Permutation invariance and positional encoding
In the proposed weight generation model, as described by the function mapping in Eq. (1), the specific arrangement of the capability vectors corresponding to different decision-makers is inconsequential. This implies that permuting the order of the decision-makers or their associated capability vectors will result in a corresponding permutation of the output combination weights, while the final output and overall performance of the architecture remain unchanged. Thus, we can express this property mathematically as follows:
$a(w_{1},...w_{m},\mathbf{s}_{1},...,\mathbf{s}_{m})$
$=a(w_{\sigma(1)},...,w_{\sigma(1)},\mathbf{s}_{\sigma(1)},...,\mathbf{s}_{\sigma(m)})$ (5)
$a(f(\mathbf{c}_{1},...,\mathbf{c}_{m};\mathbf{x};\mathbf{\theta}),\mathbf{s}_{1},...,\mathbf{s}_{m})$
$=a(f(\mathbf{c}_{\sigma(1)},...,\mathbf{c}_{\sigma(m)};\mathbf{x};\mathbf{\theta}),\mathbf{s}_{\sigma(1)},...,\mathbf{s}_{\sigma(m)})$ (6)
where $\sigma(1),...,\sigma(m)$ represents a permutation of the original indices $1,...,m$. Given that both absolute and relative positions are irrelevant in this context, we do not incorporate positional encoding into the capability embeddings. Instead, we utilize a token type embedding to differentiate between the capability embeddings and the task embedding. This approach ensures that the model maintains its performance characteristics regardless of the arrangement of the decision-makers.
@@ 5 A Learning-free Baseline for Multi-class Classification
In the case of classification problem, some existing works assume that a human expert can predict the correct label for a subset of classes, while making random selections for other classes (Mozannar and Sontag, 2020). However, in real world, an educated human expert does have the ability of doing some level of correct choice. For example, it is not likely that one classifies an animal as a vehicle, while sometime one may “not sure” if a photo is about one person or another. Therefore, we can assume that a human expert has one or multiple expertise categories, in which they can make close to 100% correct decisions. For other categories, they can make correct decision with a limited probability $p$, and make random choice in other cases with probability $1-p$.
@@ 5.2.1 Building the combined output vector
    $b_{y}$ $=\alpha a^{1}_{y}+(1-\alpha)$ (7)
    $b_{j}$ $=\alpha a^{1}_{j}$ (8)
@@
    $b_{y}$ $=\alpha a^{1}_{y}+(1-\alpha),\quad P=p+\frac{1}{n-K}(1-p)$ (9)
    $b_{y}$ $=\alpha a^{1}_{y},\quad P=\frac{n-K-1}{n-K}(1-p)$ (10)
    $b_{j}$ $=\alpha a^{1}_{j},\quad P=p+\frac{n-K-1}{n-K}(1-p)$ (11)
    $b_{j}$ $=\alpha a^{1}_{j}+(1-\alpha),\quad P=\frac{1}{n-K}(1-p)$ (12)
@@ 5.2.2 Model output approximation
To model the noise introduced to the one-hot representation, we employ a normal distribution $\delta_{j}\sim N(0,\sigma^{2})$, assuming that the noise terms are independent and identically distributed (i.i.d.). The process of deriving the predicted label from the model output under these assumptions is illustrated in Equation (13):
$[a_{1},\ldots,a_{y},\ldots,a_{n}]\stackrel{{\scriptstyle\text{map}}}{{\rightarrow}}[\delta_{1},\ldots,1+\delta_{y},\ldots,\delta_{n}]$
$\stackrel{{\scriptstyle\text{denoise}}}{{\rightarrow}}[0,\ldots,1,\ldots,0]$ (13)
@@ 5.2.3 Searching for the optimal combination weights
    $b_{y}$ $=\alpha(\delta_{y}+1)+(1-\alpha)=\alpha\delta_{y}+1,$ (14)
    $b_{j}$ $=\alpha\delta_{j}$ (15)
@@
    $b_{y}$ $=\alpha(\delta_{y}+1),$ (16)
    $b_{j}$ $=\alpha\delta_{j}+(1-\alpha)\quad\text{or}\quad b_{j}=\alpha\delta_{j}$ (17)
@@
$\begin{split}P_{1}(b_{y}>b_{j})=&P(y\in E,b_{y}>\max\{b_{j}\}_{j\neq y})+P(y\notin E,b_{y}>\max\{b_{j}\}_{j\neq y})\\ =&P(y\in E)P(b_{y}>\max\{b_{j}\}_{j\neq y}|y\in E)+P(y\notin E)P(b_{y}>\max\{b_{j}\}_{j\neq y}|y\notin E)\\ =&\frac{K}{n}P(\delta_{y}-\max\{\delta_{j}\}>-\frac{1}{\alpha})+\frac{n-K}{n}(P(\delta_{y}-\delta_{{j}_{1}}>\frac{1-2\alpha}{\alpha},\\ &\delta_{y}-\max\{\delta_{j}\}>-1)P_{b}+P(\delta_{y}-\max\{\delta_{j}\}>-\frac{1}{\alpha})P_{a})\\ =&\frac{K}{n}\prod_{j\neq y}P(\delta_{y}-\delta_{j}>-\frac{1}{\alpha})+\frac{n-K}{n}(P_{b}P(\delta_{y}-\delta_{{j}_{1}}>\frac{1-2\alpha}{\alpha})\\ &\prod_{j\neq y,j_{1}}P(\delta_{y}-\delta_{j}>-1)+P_{a}\prod_{j\neq y}P(\delta_{y}-\delta_{j}>-\frac{1}{\alpha}))\end{split}$ (18)
@@
$\begin{split}F(\alpha)=&\frac{K}{n}(\int^{\infty}_{-\frac{1}{\alpha}}\exp^{-\frac{x^{2}}{2\sigma_{D}^{2}}}dx)^{n-1}\\ &+\frac{n-K}{n}(P_{b}(\int^{\infty}_{\frac{1-2\alpha}{\alpha}}\exp^{-\frac{x^{2}}{2\sigma_{D}^{2}}}dx)(\int^{\infty}_{-1}\exp^{-\frac{x^{2}}{2\sigma_{D}^{2}}}dx)^{n-2}\\ &+P_{a}(\int^{\infty}_{-\frac{1}{\alpha}}\exp^{-\frac{x^{2}}{2\sigma_{D}^{2}}}dx)^{n-1})-(\int^{\infty}_{-1}\exp^{-\frac{x^{2}}{2\sigma_{D}^{2}}}dx)^{n-1}\\ &=\frac{K}{n}(\Phi(\frac{1}{\alpha\sigma_{D}}))^{n-1}+\frac{n-K}{n}(P_{b}(\Phi(\frac{2\alpha-1}{\alpha\sigma_{D}})\Phi(1/\sigma_{D})^{n-2}\\ &+P_{a}(\Phi(\frac{1}{\alpha\sigma_{D}}))^{n-1})-(\Phi(\frac{1}{\sigma_{D}}))^{n-1}\\ >&0\end{split}$ (20)
@@ 6 Experiments
We focus on two image classification datasets and one sentiment classification dataset for collaborative decision-making, including CIFAR10, CIFAR100 and Hatespeech. CIFAR-10 (Krizhevsky et al., 2009) comprises 60,000 color images, each measuring 32x32 pixels. These images are categorized into 10 distinct classes, with each class containing 6,000 images. The dataset is divided into 50,000 training images and 10,000 test images. CIFAR-100 consists of 60,000 32x32 color images spread across 100 classes, which are further organized into 20 super-classes (Krizhevsky et al., 2009). Each class contains 600 images, with 5 classes per super-class. Each image is annotated with a “fine” label indicating its specific class and a “coarse” label representing its super-class. Hatespeech dataset consist of 24,783 tweets annotated as hate speech, offensive language or neither (Davidson et al., 2017). For our experiments with CIFAR-10 and CIFAR-100, we utilize the first 45,000 instances from the training set as the training data, reserving the remaining 5,000 instances for validation, while the original test set is employed as the test data. In the case of the Hatespeech dataset, we use the first 80% of the data for training, the following 10% for validation, and the rest 10% for testing.
@@ 6.3.1 Fixed decision makers
At first, we consider a simple scenario. Assume we have two human experts. The first is familiar with animals, and the second is familiar with vehicles. However, we do not know the expertise of them before comparing their decision with the ground-truth labels. We assume that for CIFAR10 dataset, the first human expert can make correct classification on the images of birds, cats, horse, dog deer and frog, while the second human expert can make correct classification on the images of airplane, automobile, ship and truck. As there are only two decision makers, we can use two different capability vector to represent their decision capacity on image classification. We set $\mathbf{c}_{1}=(0,1)$ and $\mathbf{c}_{2}=(1,0)$, and map these two capability vectors to capability embedding vectors with linear transformations $c^{\prime}_{1}=\mathbf{c}_{1}\mathbf{W}_{c}$, $c^{\prime}_{2}=\mathbf{c}_{2}\mathbf{W}_{c}$, where $c^{\prime}_{1},c^{\prime}2\in\mathbb{R}^{1\times d}$, $\mathbf{W}_{c}\in\mathbb{R}^{2\times d}$, and $d$ is the input embedding dimension. Meanwhile, we apply a pre-trained ResNet-18 to extract the feature of images (He et al., 2016), where the top-level representation has a length of 512. Thus we set the input embedding dimension to be 512. We use a 4-layer 8-head transformer encoder with hidden size equal to 512. We use a transformer encoder as the weights generation model, and apply the segment embeddings to distinguish the positions of capability vectors and task embeddings. An Adam optimizer with a learning rate equals to 1e-5 is applied for optimization, while the batch size is set to be 64. We compare three settings of the feature extractor: a. No task embedding: We only apply capability vectors without using task embedding to determine the combination weights of decision makers, which results in the same combination weights for all task instance; b. Fixed task embedding: We use a finetuned Resnet-18 on CIFAR10 with fixed model parameters; c. Full version: We initialize the model with a Resnet-18 finetuned on CIFAR10, while the parameter can be further updated during the end-to-end training on collaborative decision making tasks. The results are given in Table 2. It shows that the proposed collaboration mechanism with tunable image representation extractor achieve the optimal classification accuracy. Given this finding, we use flexible feature extractors in the following experiments.

### v2

@@ 3.2 Learning mechanism of capability vectors
The learning mechanism for capability vectors involves updating these vectors based on collaborative decision-making performance. As is shown in Figure 2, each decision maker, whether a human expert or an AI model, is initially assigned a unique one-hot vector. This vector is then transformed into a capability vector through a linear transformation:
$\mathbf{c}_{i}=\mathbf{e}_{i}\mathbf{W}_{c}$ (4)
Here, $\mathbf{e}_{i}$ is the one hot vector of $i$th expert, $\mathbf{W}_{c}\in\mathbb{R}^{d_{\text{vocab}}\times d_{\text{cap}}}$ is a learnable weight matrix, with $d_{\text{vocab}}$ representing the number of decision agents and $d_{\text{cap}}$ the dimensionality of the capability vectors. These capability vectors are fed into a weight generation model along with task embedding. Optimized capability vectors are crucial for determining optimal combination weights for candidate agents’ outputs. If the capability vectors misrepresent the decision agent’s actual capabilities, the weight generation model cannot assign accurate weights, leading to reduced decision accuracy.
@@
The final decision scores are computed by the weighted sum of the scores from all decision agents, using the generated weights. The capability vectors are updated through back-propagation using gradients derived from the final decision loss. The update rule for the capability matrix is:
$\mathbf{W}_{c}\leftarrow\mathbf{W}_{c}-\eta\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{c}}$ (5)
where $\eta$ is the learning rate and $\mathcal{L}$ represents the mini-batch loss function. Our learning mechanism employs random mini-batch sampling to train capability vectors, dynamically forming each batch by randomly sampling decision agents and tasks. This stochastic approach ensures the model generalizes well across different decision agents and tasks.
This mechanism resembles the training of token embeddings in language models and user embeddings in recommendation systems (Mikolov et al. 2013, He et al. 2017). However, capability vectors uniquely focus on decision-making performance across tasks. By refining capability vectors through context-based learning, we ensure they effectively capture the decision-making capabilities of both human experts and AI models.
3.3 Computational Complexity Analysis
The proposed dynamic task-aware collaboration framework exhibits the following computational complexity characteristics. The core operations are dominated by three components: (1) Capability vector learning via linear transformation of one-hot decision agent identifiers Eq. (3), (2) Weight generation through a Transformer encoder that processes capability vectors and task embeddings Eq. (1), and (3) Decision score aggregation via weighted matrix combination Eq. (2). Key complexity parameters include: $m$ (active decision agents per task), $k$ is the length of task embedding (e.g. for image classification using ResNet output feature, $k=1$), the number of candidate decision choices $n$ , embedding dimension $d$ (512 by default), the number of transformer layers $L$ (6 by default), and the number of attention heads $h$ (8 by default). The quadratic $O(m^{2})$ term in Transformer attention constitutes the primary scalability bottleneck for large $m$, though experimental results demonstrate feasibility up to $m=1000$ with standard hardware. The detailed comparison is shown in Table 1.
Table 1: Computational complexity of framework components
Component Operation Time Complexity Space Complexity
Cap. Vec. Learning One-hot to vector mapping $O((m+k)d)$ $O(d_{\text{vocab}}d)$
$\mathbf{c}_{i}=\mathbf{e}_{i}\mathbf{W}_{c}$ per batch
Weight Generation Transformer encoder $O\bigl(Lh\bigl((m+k)^{2}d+(m+k)d^{2}\bigr)\bigr)$ $O\bigl(Lh\bigl((m+k)^{2}+d^{2}\bigr)\bigr)$
$f(\mathbf{c}_{1:(m+k)},\mathbf{x};\theta)$ per instance
Decision Aggregation Weighted score fusion $O((m+k)n)$ $O((m+k)n)$
$\mathbf{s}_{f}=\mathbf{w}\mathbf{S}$ per instance
Overall System Training (end-to-end) $O\bigl(Lh(m+k)^{2}d+Lh(m+k)d^{2}$ $(m+k)n\bigr)$ $O\bigl(Lh(m+k)^{2}+d^{2}$ $(m+k)n\bigr)$
Inference (per sample) $O\bigl(Lh(m+k)^{2}d+Lh(m+k)d^{2}$ $(m+k)n\bigr)$ $O\bigl(Lh(m+k)^{2}+d^{2}\bigr)$
@@ 3.4 Permutation invariance and positional encoding
The proposed weight generation model (as described by Eq.  (1)) demonstrates permutation invariance, meaning the order of decision agents or their capability vectors doesn’t affect the final output architecture or performance. Mathematically, permuting indices in the capability vectors results in correspondingly permuted combination weights without altering overall performance:
$a(w_{1},...w_{m},\mathbf{s}_{1},...,\mathbf{s}_{m})$ $=a(w_{\sigma(1)},...,w_{\sigma(1)}$
$,\mathbf{s}_{\sigma(1)},...,\mathbf{s}_{\sigma(m)})$ (6)
$a(f(\mathbf{c}_{1},...,\mathbf{c}_{m};\mathbf{x};\mathbf{\theta}),\mathbf{s}_{1},...,\mathbf{s}_{m})$ $=a(f(\mathbf{c}_{\sigma(1)},...,\mathbf{c}_{\sigma(m)};$
$\mathbf{x};\mathbf{\theta}),\mathbf{s}_{\sigma(1)},...,\mathbf{s}_{\sigma(m)})$ (7)
where $\sigma(1),...,\sigma(m)$ represents a permutation of the original indices 1,,m1,...,m. Given this permutation invariance, positional encoding is omitted in capability embeddings. Instead, a token type embedding differentiates capability embeddings from task embeddings, ensuring consistent model performance regardless of decision agent arrangement. This approach maintains the architecture’s effectiveness while accommodating variable decision agent orders.
@@ 4 A Learning-free Baseline for Multi-class Classification
In the case of classification problem, some existing works assume that a human expert can predict the correct label for a subset of classes, while making random selections for other classes (Mozannar and Sontag 2020). However, in real world, an educated human expert does have the ability of doing some level of correct choice. For example, it is not likely that one classifies an animal as a vehicle, while sometime one may “not sure” if a photo is about one person or another. Therefore, we can assume that a human expert has one or multiple expertise categories, in which they can make close to 100% correct decisions. For other categories, they can make correct decision with a limited probability $p$, and make random choice in other cases with probability 1p1-p.
@@ 4.2.1 Building the combined output vector
    $b_{y}$ $=\alpha a^{1}_{y}+(1-\alpha)$
    $b_{j}$ $=\alpha a^{1}_{j}$
@@
    $b_{y}$ $=\alpha a^{1}_{y}+(1-\alpha),\quad P=p+\frac{1}{n-K}(1-p)$
    $b_{y}$ $=\alpha a^{1}_{y},\quad P=\frac{n-K-1}{n-K}(1-p)$
    $b_{j}$ $=\alpha a^{1}_{j},\quad P=p+\frac{n-K-1}{n-K}(1-p)$
    $b_{j}$ $=\alpha a^{1}_{j}+(1-\alpha),\quad P=\frac{1}{n-K}(1-p)$
@@ 4.2.2 Model output approximation
To model the noise introduced to the one-hot representation, we employ a normal distribution $\delta_{j}\sim N(0,\sigma^{2})$, assuming that the noise terms are independent and identically distributed (i.i.d.). The process of deriving the predicted label from the model output under these assumptions is illustrated by:
$a_{1},\ldots,a_{y},\ldots,a_{n}]\stackrel{{\scriptstyle\text{map}}}{{\rightarrow}}[\delta_{1},\ldots,1+\delta_{y},\ldots,\delta_{n}]\stackrel{{\scriptstyle\text{denoise}}}{{\rightarrow}}[0,\ldots,1,\ldots,0]$
@@ 4.2.3 Searching for the optimal combination weights
    $b_{y}$ $=\alpha(\delta_{y}+1)+(1-\alpha)=\alpha\delta_{y}+1,$
    $b_{j}$ $=\alpha\delta_{j}$
@@
    $b_{y}$ $=\alpha(\delta_{y}+1),$
    $b_{j}$ $=\alpha\delta_{j}+(1-\alpha)\quad\text{or}\quad b_{j}=\alpha\delta_{j}$
@@
$P_{1}(b_{y}>b_{j})=$ $P(y\in E,b_{y}>\max\{b_{j}\}_{j\neq y})+P(y\notin E,b_{y}>\max\{b_{j}\}_{j\neq y})$
$=$ $P(y\in E)P(b_{y}>\max\{b_{j}\}_{j\neq y}|y\in E)+P(y\notin E)P(b_{y}>\max\{b_{j}\}_{j\neq y}|y\notin E)$
$=$ $\frac{K}{n}P(\delta_{y}-\max\{\delta_{j}\}>-\frac{1}{\alpha})+\frac{n-K}{n}(P(\delta_{y}-\delta_{{j}_{1}}>\frac{1-2\alpha}{\alpha},$
$\delta_{y}-\max\{\delta_{j}\}>-1)P_{b}+P(\delta_{y}-\max\{\delta_{j}\}>-\frac{1}{\alpha})P_{a})$
$=$ $\frac{K}{n}\prod_{j\neq y}P(\delta_{y}-\delta_{j}>-\frac{1}{\alpha})+\frac{n-K}{n}(P_{b}P(\delta_{y}-\delta_{{j}_{1}}>\frac{1-2\alpha}{\alpha})$
$\prod_{j\neq y,j_{1}}P(\delta_{y}-\delta_{j}>-1)+P_{a}\prod_{j\neq y}P(\delta_{y}-\delta_{j}>-\frac{1}{\alpha})),$ (8)
@@
$F(\alpha)=$ $\frac{K}{n}(\int^{\infty}_{-\frac{1}{\alpha}}\exp^{-\frac{x^{2}}{2\sigma_{D}^{2}}}dx)^{n-1}$
$+\frac{n-K}{n}(P_{b}(\int^{\infty}_{\frac{1-2\alpha}{\alpha}}\exp^{-\frac{x^{2}}{2\sigma_{D}^{2}}}dx)(\int^{\infty}_{-1}\exp^{-\frac{x^{2}}{2\sigma_{D}^{2}}}dx)^{n-2}$
$+P_{a}(\int^{\infty}_{-\frac{1}{\alpha}}\exp^{-\frac{x^{2}}{2\sigma_{D}^{2}}}dx)^{n-1})-(\int^{\infty}_{-1}\exp^{-\frac{x^{2}}{2\sigma_{D}^{2}}}dx)^{n-1}$
$=\frac{K}{n}(\Phi(\frac{1}{\alpha\sigma_{D}}))^{n-1}+\frac{n-K}{n}(P_{b}(\Phi(\frac{2\alpha-1}{\alpha\sigma_{D}})\Phi(1/\sigma_{D})^{n-2}$
$+P_{a}(\Phi(\frac{1}{\alpha\sigma_{D}}))^{n-1})-(\Phi(\frac{1}{\sigma_{D}}))^{n-1}$
$>$ $0$ (10)
@@ 5 Experiments
We evaluate our method on four image classification datasets and one sentiment classification dataset for collaborative decision-making, including CIFAR10, CIFAR-10H, CIFAR100, GalaxyZoo, and Hatespeech. CIFAR-10 (Krizhevsky et al. 2009) comprises 60,000 color images, each measuring 32x32 pixels. These images are categorized into 10 distinct classes, with each class containing 6,000 images. The dataset is divided into 50,000 training images and 10,000 test images. CIFAR-100 consists of 60,000 32x32 color images spread across 100 classes, which are further organized into 20 super-classes (Krizhevsky et al. 2009). Each class contains 600 images, with 5 classes per super-class. Each image is annotated with a “fine” label indicating its specific class and a “coarse” label representing its super-class. Hatespeech dataset consist of 24,783 tweets annotated as hate speech, offensive language or neither (Davidson et al. 2017). For our experiments with CIFAR-10 and CIFAR-100, we utilize the first 45,000 instances from the training set as the training data, reserving the remaining 5,000 instances for validation, while the original test set is employed as the test data. In the case of the Hatespeech dataset, we use the first 80% of the data for training, the following 10% for validation, and the rest 10% for testing. In addition, CIFAR-10H and GalaxyZoo are datasets with real human labels. CIFAR-10H (Peterson et al. 2019) is a human-annotated variant of the CIFAR-10 test set, where each of the 10,000 images has been labeled by multiple human annotators. GalaxyZoo is a crowd-sourced astronomical dataset for galaxy morphology classification (Lintott et al. 2008). The detailed pre-processing procedure of these two datasets is given in Section 5.4.4.
@@ 5.3.1 Fixed decision makers
At first, we consider a simple scenario. Assume we have two human experts. The first is familiar with animals, and the second is familiar with vehicles. However, we do not know the expertise of them before comparing their decision with the ground-truth labels. We assume that for CIFAR10 dataset, the first human expert can make correct classification on the images of birds, cats, horse, dog deer and frog, while the second human expert can make correct classification on the images of airplane, automobile, ship and truck. As there are only two decision makers, we can use two different capability vector to represent their decision capacity on image classification. We set $\mathbf{c}_{1}=(0,1)$ and $\mathbf{c}_{2}=(1,0)$, and map these two capability vectors to capability embedding vectors with linear transformations $c^{\prime}_{1}=\mathbf{c}_{1}\mathbf{W}_{c}$, $c^{\prime}_{2}=\mathbf{c}_{2}\mathbf{W}_{c}$, where $c^{\prime}_{1},c^{\prime}2\in\mathbb{R}^{1\times d}$, $\mathbf{W}_{c}\in\mathbb{R}^{2\times d}$, and $d$ is the input embedding dimension. Meanwhile, we apply a pre-trained ResNet-18 to extract the feature of images (He et al. 2016), where the top-level representation has a length of 512. Thus we set the input embedding dimension to be 512. We use a 4-layer 8-head transformer encoder with hidden size equal to 512. We use a transformer encoder as the weights generation model, and apply the segment embeddings to distinguish the positions of capability vectors and task embeddings. An Adam optimizer with a learning rate equals to 1e-5 is applied for optimization, while the batch size is set to be 64. We compare three settings of the feature extractor: a. No task embedding: We only apply capability vectors without using task embedding to determine the combination weights of decision makers, which results in the same combination weights for all task instance; b. Fixed task embedding: We use a finetuned Resnet-18 on CIFAR10 with fixed model parameters; c. Full version: We initialize the model with a Resnet-18 finetuned on CIFAR10, while the parameter can be further updated during the end-to-end training on collaborative decision making tasks. The results are given in Table 3. It shows that the proposed collaboration mechanism with tunable image representation extractor achieve the optimal classification accuracy. Given this finding, we use flexible feature extractors in the following experiments.